### v1

# scenarioparampadding=8pt, margin-top=3pt, margin-bottom=3pt, border-style=none, background-color=quotebackground

Abstract.
In the mental health domain, Large Language Models (LLMs) offer promising new opportunities, though their inherent complexity and low controllability have raised questions about their suitability in clinical settings. We present MindfulDiary, a mobile journaling app incorporating an LLM to help psychiatric patients document daily experiences through conversation. Designed in collaboration with mental health professionals (MHPs), MindfulDiary takes a state-based approach to safely comply with the experts’ guidelines while carrying on free-form conversations. Through a four-week field study involving 28 patients with major depressive disorder and five psychiatrists, we found that MindfulDiary supported patients in consistently enriching their daily records and helped psychiatrists better empathize with their patients through an understanding of their thoughts and daily contexts. Drawing on these findings, we discuss the implications of leveraging LLMs in the mental health domain, bridging the technical feasibility and their integration into clinical settings.
journaling, chatbot, mental health, clinical setting, psychiatric patient, large language models
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06submissionid: 3044ccs: Human-centered computing Empirical studies in HCIccs: Human-centered computing Natural language interfaces
Figure 1. MindfulDiary consists of two components: the patient interface and the expert interface: (1) Patient interface aids users in daily journaling, offering prompt questions through conversations and crafting summaries in a diary-like format. (2) Clinician Dashboard features a dashboard that visualizes data from patient records, facilitating its use during consultations. (Refer to our supplementary video for the user interaction.)
1. Introduction
Journals serve as a written record of an individual’s past events, thoughts, and feelings, allowing genuine expression  (Travers, 2011; Ullrich and Lutgendorf, 2002). Prior work has shown the advantages of journaling in clinical mental health contexts, as journals frequently capture patients’ daily experiences, symptoms, and other contextual data that are challenging to gather during brief hospital visits (Figueiredo et al., 2020; Zhu et al., 2016). Furthermore, these patient journals can enhance mental health professionals’ (MHPs) comprehension of their patients’ conditions, leading to improved treatment quality (Wu et al., 2020). However, writing about one’s past feelings and thoughts can be a complex process because people differ in their ability to understand, identify, and verbalize their emotions (Salovey and Mayer, 1990). In addition, patient under psychotherapy struggle with constructing a narrative and understanding their past (Donnelly and Murray, 1991; Pennebaker and Seagal, 1999).
Researchers in the field of Human-Computer Interaction (HCI) have shown that chatbots can help individuals articulate and share their daily experiences. For instance, chatbots to elicit people’s self-disclosure can ease the process of emotional expression by providing a safe and supportive environment for individuals to share their experiences and emotions (Park et al., 2021; Lee et al., 2020b, a; Cameron et al., 2019). Furthermore, a machine’s inherent trait of not showing fatigue can make people more confident to share their stories truthfully and comfortably (Kim et al., 2018; Park et al., 2021).
The recent advancements in large language models (LLMs) have added to the understanding and generation of natural conversations. Unlike traditional chatbot (Abd-alrazaq et al., 2019) were unable to support conversation outside pre-defined context (Wang et al., 2021; Dsouza et al., 2019), chatbot driven by LLMs can offer dynamic responses on diverse topics (Bae et al., 2022; Roller et al., 2021; Brown et al., 2020). However, care needs to be taken when leveraging LLMs, particularly in mental health contexts, as the inherent challenge of LLMs in uncertainty in control could result in generating unintended or inaccurate responses (Farhat, 2023; Wang et al., 2021; Korngiebel and Mooney, 2021).
In this work, we present MindfulDiary, a mobile conversational AI to support patients in journaling their daily experiences and thoughts through conversations. MindfulDiary incorporates LLMs to generate a response, prompting patients differently according to the conversational phase. The conversation records are automatically summarized and presented on a clinician dashboard so MHPs can obtain insights about the patient. As a multi-disciplinary research team, which included HCI researchers, AI engineers, and psychiatrists, we iteratively designed MindfulDiary and conducted a four-week field study with 28 psychiatric patients diagnosed with major depressive disorder (MDD). Through this study, we found that the versatility, narrative-building capability, and diverse perspectives provided by MindfulDiary assisted patients in consistently enriching their daily records. Furthermore, MindfulDiary supported patients in overcoming the challenges of detailed record-keeping and expression, often hindered by feelings of apathy and cognitive burdens. Notably, the enriched records from MindfulDiary provided MHPs with deeper insights, enhancing their understanding and empathy toward their patients. Based on the findings, we discuss the potential of LLMs’ conversational abilities to aid in patient record-keeping within clinical mental health settings and suggestions for their responsible integration into these environments.
Our work contribute the following:
  1. (1)
    We present MindfulDiary, an LLM-driven journal designed to document psychiatric patients’ daily experiences through naturalistic conversations, designed in collaboration with MHPs.
  2. (2)
    From a four-week field study involving 28 patients and five psychiatrists, we provide empirical findings on how MindfulDiary supported patients in keeping their daily logs and assisted psychiatrists in monitoring and comprehending patient states.
  3. (3)
    Drawing from the design process of MindfulDiary and the findings of our deployment study, we outline key considerations and share lessons learned for the design and implementation of LLM’s conversational capability in clinical mental health settings.
2. Related Work
In this section, we cover related work in two parts: (1) journal as a patient-generated health data, and (2) conversational agent for mental health.
2.1. Journal as a Patient-generated Health Data in Clinical Setting
PGHD—defined as “health-related data, such as health history, symptoms, biometric readings, treatment history, lifestyle choices, and other pertinent details, created, recorded, or gathered by patients” (Shapiro et al., 2012)—has increasingly become an essential tool in clinical settings to capture authentic, real-time insights into patients‘ health. Studies have shown that PGHD can enhance communication between patients and MHPs and offer contextual information about patients, thereby heightening MHPs’ awareness of patient health outside regular clinical visits (Lordon et al., 2020; Cohen et al., 2016; Nundy et al., 2014).
Within the mental health domain, PGHD range from structured mental health assessments (e.g., anxiety, depression) to more unstructured data, including mood-related symptoms and social interactions tracking (e.g., social media use, number of calls)(Austin et al., 2020; Demiris et al., 2019; Nghiem et al., 2023). Our study places particular emphasis on journals as a PGHD method in clinical mental health settings. Journaling refers to “the practice of writing down one‘s symptoms and other information related to one‘s daily life to discuss them during clinical appointments” (Figueiredo et al., 2020; Zhu et al., 2016). Using a journal as a PGHD can particularly be useful in mental health contexts as it can offer rich, self-documented insights, which could improve MHPs‘ understanding of their patients (Wu et al., 2020).
However, while the benefits of journaling are clear, the act itself is not always straightforward. Many people struggle with starting their entries, sticking to consistent journaling routines, and structuring their reflections (Travers, 2011). Further, writing about emotions and past experiences can be intricate, as individuals vary in their capacity to recognize, interpret, and articulate their feelings (Salovey and Mayer, 1990; Lane et al., 1990; Luminet et al., 2004). For some, especially in psychotherapy, crafting a narrative that describes one’s life journey can be a challenging process (Donnelly and Murray, 1991; Pennebaker and Seagal, 1999).
While previous research has extensively established the significance of employing journals as PGHD within clinical contexts, it is essential to acknowledge the inherent complexities tied to the journaling process. This becomes especially critical when contemplating its application within the specific demographic undergoing psychotherapy. With this background, our research explores the potential of LLMs in assisting the journaling process within the clinical mental health setting.
2.2. Conversational Agent for Mental Health
The field of artificial intelligence (AI) has proposed some significant innovations in medical settings, such as aiding clinical decision-making and diagnosis (Magrabi et al., 2019; Beam and Kohane, 2016). However, AIs have rarely been integrated into mental health contexts, partly because the field has relied on nuanced techniques and subjective data, such as patient relationships and emotion observation (Graham et al., 2019; Gabbard and Crisp-Han, 2017). Recently, there has been a growing technological presence of Natural Language Processing (NLP) applications for supporting mental health, as NLP’s capability for understanding human language holds a significant promise in this domain (Calvo et al., 2017; Demner-Fushman et al., 2009).
Conversational agents, also known as chatbots, have particularly stood out in the mental health domain (Abd-Alrazaq et al., 2021; Abd-alrazaq et al., 2019). Studies have demonstrated the potential of chatbots in facilitating different types of therapy, such as cognitive behavioral therapy (Fitzpatrick et al., 2017; Inkster et al., 2018; Fulmer et al., 2018), expressive writing (Park et al., 2021), behavioral reinforcement (Inkster et al., 2018), and solution-focused therapy (Fulmer et al., 2018). Prior work has also shown that chatbots could ease the burden of disclosing sensitive information. Studies indicated that individuals may feel more comfortable communicating with chatbots because of the social stigma involved in communicating with human beings (Lee et al., 2020b; Park et al., 2021; Lee et al., 2020a). Furthermore, these approaches can help overcome temporal and spatial constraints, offering mental health support that is accessible at any time and anywhere (Cameron et al., 2019).
Recently, chatbots powered by LLMs have gained attention (Wang et al., 2021). Trained on vast amounts of text data, they demonstrate exceptional performance and versatility in natural open-domain conversation, surpassing previous NLP models in understanding intricate language patterns (Radford et al., 2019; Tanana et al., 2021). Yet, many chatbots used for mental health therapy still employ a retrieval-based approach (Abd-alrazaq et al., 2019; Abd-Alrazaq et al., 2019). The basic principle of retrieval-based chatbots is to search for an appropriate response from a database based on input from users. Since the responses and expected flow that the system can provide are already hard-coded in the database, chatbots exhibit consistent and predictable behavior (Dsouza et al., 2019). However, if a user gives input that does not exist in the database, the chatbot may fail to respond. Thus, these systems cannot flexibly respond to unforeseen situations. In contrast, a generative approach driven by LLM has made it possible to develop chatbots that can produce human-like and contextualized outputs for even unforeseen input, enabling open-domain dialogues (Bae et al., 2022).
However, LLMs, employing a neural network known as a “transformer” to generate outputs, come with inherent challenges tied to their architecture (Vaswani et al., 2017). One key issue is the explainability of the model output: it’s challenging to discern how this ‘black box’ model interprets a given input prompt. As a result, designers struggle to predict both the LLM’s understanding of the input and the subsequent messages it might produce (Liu et al., 2023). For instance, a chatbot leveraging GPT-2 for mental health therapy occasionally generated non-word sentences or produced more negative outputs than positive ones (Wang et al., 2021). Replika, an LLM-based application intended for mental well-being, has occasionally displayed harmful content and exhibited inconsistent conversational styles, undermining its role as a long-term companion (Ma et al., 2023).
Considering the inherent challenges of LLM and the sensitivity of the mental health domain, we must reflect carefully on how and to what extent the conversational capabilities of these models should be utilized. In this study, we emphasize the need to consider the opportunities and limitations of LLMs, as well as important aspects of the mental health domain, including safety, sensitivity, and experts’ empirical knowledge. With this in mind, we designed MindfulDiary’s state-based chatbot using both rule-based and generation-based approaches to safe compliance with experts’ guidelines while facilitating free-form conversations.
3. Formative Study: Focus Group Interview
To inform the design of MindfulDiary, we first conducted a Focus Group Interview (FGI) with MHPs. We recruited six MHPs (E1–6; two males and four females)—four clinical psychologists and two psychiatrists—from a local university hospital in Korea. We held two one-hour remote sessions via Zoom. In the FGI, we first provided an overview of language model technologies and LLM’s natural language understanding and generation capabilities. We also described the current challenges of LLMs, such as potential errors and hallucinations. We then asked participants a focused set of questions on (1) the challenges MHPs currently face during patient treatment and counseling sessions and (2) their expectations and envisioned opportunities of LLMs’ role in clinical mental health settings. The session was video recorded and later transcribed. We open-coded the transcripts to identify emerging themes. In the following, we cover the findings from the FGI.
3.1. Challenges in Eliciting Responses from Patients with Depression
Participants indicated that eliciting disclosure from patients’ inner thoughts during a limited consultation time requires significant effort. Many patients with depression experience difficulty describing and expressing their feelings and thoughts to providers due to a sense of apathy, which is a common psychiatric symptom involved in Major Depressive Disorder: “In the consultation room, even if they sit like this, they often just remain silent for a long time.” (E5) Thus, providers often end up spending a substantial amount of time asking standardized and repetitive questions about mood, sleep, and major events to understand patients’ current states.
Participants also noted that they had their patients engaged in paper-based diary writing programs but most demonstrated low participation rates and low engagement: “We tried the diary program on paper(in the inpatient ward), and several patients did write. What we saw was quite trivial, like, ‘I just felt bad today.’ But we learned there were significant events upon consultation, like having a big argument with other patients, which they did not record. Because patients with depression, or those who have had suicidal or self-harming incidents, often have a dulled state in expressing their emotions or feel apathetic, they tend to find such expressions very difficult.” (E3)
Participants envisioned that language models have significant potential given their natural and flexible conversational abilities. They suggested that this ability could help patients explore their daily lives and engage in deep reflection, making the daily logging process more engaging by providing a sense of companionship.
3.2. The Meaning of Daily Records as Perceived by MHPs
Our participants emphasized that patient records about their daily experiences can serve as crucial data for MHPs to understand them. E5 envisioned that such records could be utilized to understand a patient’s automatic thoughts and reaction mechanisms concerning specific events. For instance, in a doctor’s office, a patient with social phobia might briefly say, “I feel scared in crowded places,” and the conversation might not delve much deeper. However, the true depth of this fear can be more accurately and vividly captured if the patient records their feelings immediately after exposure to a crowded place or even later that night.
Participants imagined that language model outputs about patient journaling patterns could provide valuable insights about what counseling styles they should use for individual patients: “Observing how they respond to certain inputs can be valuable data. Some patients feel comforted by mere reassurance, while others prefer more direct, pointed feedback.”
From the FGI, we refined the initial concept of the MindfulDiary program. We leveraged the conversational abilities of LLMs to help patients document their daily experiences between clinical visits. MHPs had access to the collected data to inform their clinical decision-making. Furthermore, both MHPs and the research team concur that LLMs should not act solely as the primary intervention due to their inherent limitations but should function as supportive tools for clinical consultations. The subsequent section outlines the design and development process of our system.
Figure 2. Main screens of the MindfulDiary app. (a) The main screen, (b) the journaling screen, (c) the summary screen shown when the user submitted the journal dialogue, and (d) the review screen displaying the user’s past journal.
[…]
4. MindfulDiary
Informed by the findings from FGI with MHPs, we designed and developed MindfulDiary, which consists of two main components: (1) a patient mobile app for daily record-keeping and (2) a clinician dashboard that allows professionals to access and use these daily records in a clinical setting (See Figure 1). Below, we present a fictional usage scenario to demonstrate how the system works.
{longfbox}[scenarioparam]Jane, diagnosed with chronic anxiety, frequently grapples with panic attacks. To keep track of her daily experiences, her psychiatrist recommends trying MindfulDiary as part of her treatment plan.
Every evening, Jane records her daily activities, emotions, and thoughts using the MindfulDiary app. The AI leads the conversation by asking prompt questions about the day of Jane. Additionally, the system transforms Jane’s inputs into a journal-style format, in which she can organize thoughts and reflect on them later. She can easily read the summarized content whenever she wishes to look back at past events or thoughts.
Three weeks later, during a consultation, her psychiatrist uses the expert interface of MindfulDiary to review a data-driven summary of Jane’s entries. The data helped the psychiatrist identify patterns that Jane’s anxiety often spikes during her work commute. Based on this insight, the psychiatrist refines advice and introduces specific coping strategies, fostering a more personalized approach to care.
4.1. MindfuleDiary App
The MindfulDiary app for patients aims to support people who might have difficulty journaling due to apathy and cognitive load through naturalistic conversation driven by an LLM. The app consists of a home screen containing an introduction and guide to the program (Figure 2), a journal writing screen (Figure 2, 2), and a screen to review the diary entries (Figure 2).
4.1.1. Journaling User Interface
Figure 3. Use flow of MindfulDiary’s journaling session: (1) Pre-Journaling Assessment: Users undergo a mental health survey using the modified PHQ-9 (Kroenke et al., 2001) before using MindfulDiary; (2) Users converse with MindfulDiary, documenting their day; (3) Summary Presentation: After three turns, MindfulDiary presents a diary-styled summary of the conversation so far, which can also be edited by the user. Users can continue the conversation as they want. (4) Session Closure: Once all processes are completed, MindfulDiary displays today’s mental health and diary content, concluding the journaling session.
Figure 3 illustrates the overall use flow of the journaling session, which begins with a Pre-Journaling Assessment (Figure 3-1⃝) that asks to fill out a questionnaire for mental health. The questionnaire comprised the modified PHQ-9 (Kroenke et al., 2001) and a custom open-ended question inquiring about recent attempts of self-harm or suicide. This assessment prevents users who provided any clues of suicidal or self-harm from journaling on the same day. (We cover this feature in detail in Section 5.4 Ethical Consideration.)
On the next screen, the user converses with MindfulDiary, documenting the events of the day (See Figure 2). After three turns, MindfulDiary provides a summary of the conversation as an essay. Users can edit this automatically generated summary any time. When the user ends the session by pressing the end button (Figure 2, bottom), MindfulDiary displays daily mental health insights alongside the diary content on the summary screen (See Figure 2). Users can also leave a reflection message there. Lastly, users can browse their past records in the Diary Review menu (See Figure 2).
4.1.2. Conversation Design
We designed the chatbot’s conversational behavior based on insights from psychiatry literature (Othmer and Othmer, 2002a, b), which covers foundational techniques and considerations for conducting clinical interviews. We also incorporated the hands-on clinical experiences of practicing psychiatrists.
As a result, we designed the conversation of a journaling session to follow a sequence of three stages: Rapport building, Exploration, and Wrap-up. The Rapport Building state is an ice-breaker, centered on casual exchanges about a user’s day. In this state, the assistant also shares bits of information to encourage users’ openness. As we progress to the Exploration state, the emphasis shifts to a comprehensive understanding of the user’s daily events, feelings, and thoughts, facilitated by a mix of open-ended and closed-ended queries that ensure users remain engaged and in control of the dialogue. The conversation then transitions to the Wrap-up, emphasizing completion and ensuring users have fully voiced their experiences while the system remains empathetic and receptive to any lingering topics.
Besides the three main stages, we also incorporated the Sensitive Topic state that handles the most sensitive subjects, such as self-harm or suicidal ideation. When this state is triggered, psychiatrists receive instant notifications. This allows them to oversee the conversation in real-time and step in to assist the patient if necessary. Here, the system begins by empathizing with the user, recognizing their struggles, and offering a reassuring message. Following this, the system gently probes the depth of their suicidal or self-harm thoughts. If the user expresses intense or specific plans related to self-harm or suicide, the system urges them to seek prompt assistance, either at a hospital emergency room or via the local helpline.
4.1.3. Conversational Pipeline
Figure 4. Structure of MindfulDiary’s conversational pipeline. (1) Users respond to MindfulDiary’s messages. (2) The recent turn count, current state, and whole user-MindfulDiary dialogue are fed into Dialogue Analyzer. (3) Using the output of Dialogue Analyzer, a designated state prompt, a summary of dialogue (containing overall dialogue context), and the latest three conversation turns are fed into Response Generator. The resulting response is then displayed to the user. Both the Dialogue Analyzer and Response Generator operate based on the GPT-4 LLM.
Lengthy and complex input prompts for LLMs are known to cause poor task performance (Brown et al., 2020) by partly omitting latent concepts (Wu et al., 2022). To steer the LLM to comply with the conversational design we intended diligently, we designed MindfulDiary’s dialogue system as a state machine. Each conversation stage is carried on with a dedicated input prompt, which is more succinct and clear than a single master prompt containing instructions for all stages. Figure 4 illustrates our conversation pipeline that runs each time a new user message is received. The pipeline incorporates two LLM-driven modules, a dialogue analyzer and the response generator. The dialogue analyzer handles the stage transition, returning the stage suggestion—whether to stay or move to a new stage—and a summary paragraph of the current dialogue from the current dialogue information. The system then formulates an LLM prompt, combining a prompt for the current stage, the dialogue summary, and the recent six messages. Receiving the prompt as an input, the response generator generates an AI message.
4.2. Pilot Evaluation
To ensure that MindfulDiary is reliable and safe for conversing with psychiatric patients, we underwent multiple rounds of pilot evaluation. First, we invited five psychiatrists and three clinical psychologists to test the conversational pipeline. The experts provided feedback on the instructions in the model prompts, focusing on their clinical relevance and the embedded terminology and strategies. Then, the experts inspected the chatbot’s behavior by chatting with it while role-playing as a patient persona. In particular, we examined the chatbot’s reactions to subtle implications of suicide or self-harm in user messages.
After iterating on the conversational pipeline, we conducted a pilot lab study with five patients admitted to a university hospital but about to be discharged soon. To ensure safety against risky messages from an LLM, we used a test platform where the participant’s clinician monitored the generated messages in real-time, approving them or sending better messages manually.
4.3. Clinician Dashboard
The clinician dashboard (c.f., Supplementary video) is a desktop application designed to facilitate monitoring patient’s journal entries and to provide analysis of the entries to help clinicians identify significant events, reactions, and emotions. The dashboard consists of the following components:
User Engagement. This section visualizes the participant’s overall engagement with MindfulDiary, including the number of journals written, the date and time they were written, and their length. The modified PHQ-9 scores for each session are also visualized, allowing professionals to track the user’s mental health trends using a validated tool.
Journal. This section displays the content of the journals written by patients. The information is presented in a card format, where each card offers a summary of the journal, including timestamps, total time taken to write the journal, and associated PHQ-9 score. The interaction logs between the patient and MindfulDiary are also provided in this section.
Insights. To assist professionals in browsing through the diary, this section visualizes (1) a word cloud to understand frequent terms that the participant used at a glance, (2) a summary of major events to highlight significant happenings and (3) summary of emotions to gauge the mood based on user input. When a specific period is selected for review, a comprehensive summary is generated. This summary relies on the combined power of gpt-4 and a Korean morphological analysis tool named Kiwi (Lee, 2022). Due to the limitations of language model-driven analysis, there might be occasional inaccuracies in the generated content. First-time users of this interface are alerted about possible inaccuracies. An in-interface tooltip also reminds users that the summarized outcomes might not be accurate.
4.4. Technical Implementation
MindfulDiary’s interface is developed using React, a JavaScript-based framework. The server, responsible for interfacing with the LLM and overseeing database operations, is implemented in Python. Google Firebase handles user authentication, data storage, and retrieval tasks. The conversational capabilities of MindfulDiary are powered by gpt-4, accessible through OpenAI’s API [^1] https://platform.openai.com/docs/guides/gpt/chat-completions-api. We specifically used gpt-4-0613 model. For parameter setting, we consistently set the temperature to 0.7 and both a presence penalty and frequency penalty to 0.5.
5. Field Deployment Study
Figure 5. Procedure of the Four-Week Field Deployment Study: A four-week exploration into the utilization of MindfulDiary by outpatient patients, encompassing daily use, and its integration into clinical decision-making. *Some participants did not have a follow-up visit during the experimental period
Through the four-week field deployment study with 28 patients undergoing outpatient treatment, we aimed to explore how patients and MHPs utilize MindfulDiary and what opportunities and challenges arise from its real-world use. Figure 5 illustrates the process of field deployment study.
5.1. Recruitment
We targeted outpatients from the Department of Mental Health at a University Hospital. Participants were selected based on specific criteria: (1) those who had been diagnosed with MDD and (2) those who did not exhibit heightened impulsive tendencies or harbor specific intentions towards self-harm or suicide. Eligible participants were identified through evaluations conducted by psychiatrists. Flyers and consent forms were distributed to eligible patients. For minors, the consent form process was adhered to only when they were accompanied by a guardian at the hospital.
We provided a compensation of approximately 60 USD (80,000 KRW) to participants who completed the entire four-week study process. For those who dropped out midway, we provided compensation based on the number of weeks they participated. Participants received approximately 11 USD (15,000 KRW) for every week that they were involved in (e.g., if someone participated for ten days, they would receive 11 USD). Participants were informed that consistent non-use of the system would be a criterion for study dropout. Specifically, if a participant did not engage with the MindfulDiary for three consecutive days, they would receive a notification from the researcher encouraging its use. However, continued non-use after this reminder would result in the participant being considered non-compliant with the study protocol, leading to their withdrawal from the study. This decision was made to ensure data reliability.
Through this process, a total of 36 individuals registered, of whom 8 dropped out, resulting in a final count of 28 participants who completed the 4-week field study. The majority of participants were adolescents and adults, with ages ranging from 12 to 28 years, with an average age of 17.6 ($SD=3.26$). Among the 28 participants, 11 were male, and 17 were female.
5.2. Data Collection
Patient-MindfulDiary Interaction Data. With the consent of the participants, we collected all data from their interactions with the MindfulDiary along with the raw input content and outputs from the language model.
Interviews with Patients. We conducted interviews with each patient to delve deeply into their experiences and learn how they used MindfulDiary on a daily basis. At the end of the second and fourth weeks, we conducted 15-minute debriefing interviews with participants. Considering the characteristics of depression patients, who may struggle to focus for long periods of time, the interview session was divided into two shorter sessions.
Interviews with Psychiatrists. We interviewed psychiatrists who treated the patient participants to understand how they might use the system’s data in actual clinical settings. We further gathered feedback from the psychiatrists on the opportunities and limitations of MindfulDiary, as well as suggestions for improvements.
5.3. Analysis
To explore participants’ usage patterns with MindfulDiary, we first conducted a descriptive statistics analysis. To determine any shifts in participants’ adherence over time, we examined weekly writing frequencies using a one-way repeated measures ANOVA (RM-ANOVA) with Greenhouse-Geisser correction. To gain a deeper qualitative understanding of the messages produced by MindfulDiary and interviews with patients and psychiatrists, we used open coding paired with thematic analysis (Braun and Clarke, 2006). All interviews were audio-recorded and later transcribed. The first author open-coded the interview transcripts along with the interaction log data through multiple rounds of iteration. The research team then identified patterns through discussion to formulate overarching themes.
5.4. Ethical Considerations
Our study design obtained the required approval from the Institutional Review Board (IRB) of the affiliated hospital.
Conducting this study, we are fully aware of the inherent risks associated with our research, particularly given the characteristics of participants diagnosed with MDD. To mitigate the risks, we first carefully screened participants, relying on evaluations conducted by psychiatrists. Individuals displaying heightened impulsive tendencies or harboring specific intentions towards self-harm or suicide were excluded from the study. In addition, participants were asked to take the PHQ-9 before interacting with MindfulDiary, along with an additional set of questions probing their recent attempts at self-harm or suicide. If a participant’s response to question number 9 of the PHQ-9, regarding suicidal/self-harm thoughts, scored ‘moderate or higher’ or if any recent suicide attempt was verified, the system pivoted to provide content geared towards alleviating anxiety and reducing stress rather than proceeding with the standard program. In such a case, a real-time alert was also sent to psychiatrists. Lastly, if sensitive themes frequently surfaced in a participant’s input during the study, their interactions with the program were temporarily halted. Psychiatrists subsequently re-evaluated such participants to assess the viability of their ongoing participation. During our experiment, for the case of P11, mentions of repetitive suicide and self-harm were detected. Consequently, an expert contacted the participant, the experiment was suspended for three days, and after a re-evaluation in an outpatient clinic, we resumed the program use with P11.
Further, to mitigate potential risks from the LLMs’ outputs, we embraced an iterative design methodology. The system’s interactions underwent repeated assessments to ensure it generated safe, non-harmful outputs. In addition, in the first week of each participant’s system use, all interactions between participants and MindfulDiary were observed in real time. To facilitate this process, when a participant started the session, the research team received a notification email. This notification included real-time monitoring links and reports of the survey responses that participants answered before each session. After the first week, user interactions and MindfulDiary were reviewed within a 12-hour window. During the review process, if an interaction contained sensitive content (specifically, terms pre-defined as sensitive by psychiatrists), the psychiatrists on our research team assessed the situation and contacted the affected participants if necessary.
Lastly, given that we were handling the patients’ personal and sensitive data, ensuring the secure protection and management of data was critical. Therefore, during the study, we utilized the Google Firebase authentication service to manage the user authentication process for participants. We were thus able to ensure that only authorized personnel had access to the data, and any attempts at unauthorized access could be promptly detected and managed. After the field study, all data was separated from personal identifiers to maintain anonymity.
6. Results
In this section, we report the results of the field study in four parts: (1) Journaling adherence, (2) Dialogue patterns, (3) Patients’ perspectives on MindfulDiary, and (4) MHPs’ perspectives of MindfulDiary for clinical settings.
6.1. Journaling Adherence
[…]
6.2. Dialogue Patterns
Participants and MindfulDiary exchanged a total of 4,410 messages (i.e., 2,205 pairs of the AI and participant’s messages) during the field study. Each session consisted of 10.82 messages ($SD=2.70$). The majority of exchanges between the AI and participants were carried on for an exploration of patients’ daily lives and emotions, as well as for casual conversations. In terms of the stage of the conversation, 62% (2,732 messages) of the messages were from Exploration, 30% (1220 messages) for Rapport building, and 6% (282 messages) for Sensitive topic. Only a small amount of messages were accounted for Wrapping up (62 messages) or Not-selected (14 messages).
To understand the contents that MindfulDiary generated, we delved deep into the content it generated. 72% of the AI messages took the form of questions, aiming to elicit responses about users’ daily experiences and emotions. We identified and categorized the primary strategies that MindfulDiary employed to assist patients’ journaling. There were four strategies employed by the LLM: Emotional Exploration, Activity/Behavior Exploration, In-depth Follow-up & Countermeasures, and Future Plan Exploration. For a comprehensive breakdown of these strategies, along with their descriptions and exemplar questions, refer to Table 1.
Table 1. Categorization of LLM’s strategies for generating prompts to encourage user reflection, including descriptions and exemplar questions. Strategies include Emotional Exploration, Activity Exploration, In-depth follow-up, and Future Plan Exploration with associated descriptions and examples
Category Description Example
Emotional Exploration
Messages that further inquire about the user’s emotional state, mood, and condition.
‘‘How did you feel after meeting her?’’, ‘‘Can you tell me how this situation makes you feel?’’
Activity Exploration
Messages exploring the user’s experiences or activities. These questions delve into more objective and factual content.
‘‘What kind of exercise do you do?’’, ‘‘Lunch with a friend sounds nice. What did you eat?’’
In-depth follow-up
Messages that help delve into the root causes or reasons related to mood, feeling, or specific situations.
‘‘Since when have you felt this overshadowed mood?’’, ‘‘If you feel confused, what might be the reason?’’
Future Plan Exploration
Messages inquiring about how the user reacted to or plans to respond to the events or emotions they mentioned
‘‘That situation must have been challenging. Have you considered any solutions?’’, ‘‘What have you done to alleviate the loneliness, even just a bit?’’
The average length of participants’ responses was 29.42 syllable counts, with a median of 20 (SD=35.9). This suggests a left-skewed distribution, where many participants gave shorter responses and a smaller number provided considerably longer answers, causing a high variation. The minimum response length was one character, and the maximum was 559 syllable counts. We further conducted a qualitative analysis of these responses, seeking to identify the themes present in users’ interactions with the LLM. This allowed us to understand the scope and topics of the daily records that MindfulDiary collected from the patients.
Participants interacting with MindfulDiary conveyed a range of topics (see Table 2). They described a spectrum of emotional states, from negative feelings like exhaustion and anxiety to positive sentiments of pride and joy. Events and activities were recounted, offering insights into their daily routines, such as walking during school times or decreased activity post-vacation. They also shared thoughts and beliefs, sometimes related to current events, revealing patterns linked to mental health, like feelings of exclusion and loneliness. Regarding perceived health status, comments spanned from immediate ailments, such as headaches, to long-term health challenges. Distorted perceptions about their body included content on excessive dieting. Specifically, participants frequently discussed medications, revealing not just their physical reactions but also their perceptions and behaviors toward them. Some expressed concerns over the taste, while others mentioned adverse reactions from intake, like discomfort after swallowing multiple pills at once. Lastly, the realm of relationships & interactions had participants highlighting both the challenges and supports in their interpersonal connections, revealing the significant impact these have on mental well-being, from conflicts and trust issues to moments of affirmation and encouragement.
Table 2. Summary of Participants’ Input Messages: Categories include Emotional States, Events & Activities, Thoughts & Beliefs, Perceived Health Status, and Relationships & Interactions with associated descriptions and examples
Category Description Example
Emotional states
The emotions that participants documented in their daily lives encompassed a broad spectrum, ranging from negative and depressive sentiments
“I’m so exhausted, I feel like I’m reaching my limit soon,” “I’m very worried, scared, and anxious”
Events and activities
The mentions of events, tasks, or activities they participated in or witnessed, such as exam periods or travel.
“When I attended school, I got some walking in, but after vacation, I don’t have much reason to go out, leading to a decreased activity level.”
Thoughts and beliefs
The thoughts, values, beliefs, and convictions they usually held. Including characteristic thought patterns related to mental health, such as distorted thinking
“I feel like someone is talking behind my back somewhere; they don’t like my actions and seem to exclude me”
Perceived Health Status
The physical state and health concerns, misconceptions about their health, reactions to medications, revealing their perceptions, behaviors related to drug intake
“I’ll starve and exercise to lose weight!”, “I just took my medicine, but it seems to be getting tasteless.”
Relationships & interactions
The contents about relationships with others. Difficulties stemming from interpersonal relationships, as well as the support and affirmation received from those around them
“I hated seeing my brother being happy. Forcing a cheerful tone also irked me,”
6.3. Patients’ Perspectives on MindfulDiary
Overall, participants viewed MindfulDiary as a space where they could open up and share their stories, feeling a sense of empathy from the system. Participants particularly found the dialogue-driven interactions with MindfulDiary useful. One participant, P15, mentioned, “If it was just about recording daily activities or emotions like a regular diary, it might have been less engaging, and I could’ve found it tedious or might not have persisted for long. But this felt like having a conversation with AI, which added an element of fun and kept me engaged in continuous use.” Such a dialogue-driven journaling process aided participants in maintaining consistent records and helped in forming a habit consistent with our user engagement analysis. P7 stated, “I liked chatting with the AI at first, so I kept using it. The more I used it, the more it became a habit.”
6.3.1. Broad Conversational Range: The Versatility in Documenting Diverse Interests
Our participants appreciated the LLM’s flexibility and naturalness in responding to various utterances, topics, and situations. Such broad conversational capabilities of the LLM provided participants with a space where they could document a variety of subjects tailored to individual interests and preferences. In our study, participants interacted with the LLM on diverse topics ranging from games, webcomics, novels, and movies to hobbies like making bead bracelets, allowing them to create richer and more personal records. P3 remarked, “AI systems that I have used in the past could only respond to specific words, but it is amazing how this one can respond to all sorts of things.”
6.3.2. Expanding views: Enriching Entries with Varied Perspectives
Participants also valued the diverse and new perspectives that LLM-generated responses offered, as those helped participants reflect on their daily events and emotions from various angles. This influence helped participants diverge from ruminating on depressive feelings. P12 mentioned, “ Sometimes when you note down emotions, that’s the only thought that comes to mind. Beyond that, I don’t remember much. Since MindfulDiary uses AI, my thoughts flow more easily, and I like it when it asks about different perspectives or topics.”.
6.3.3. Probing for Depth: Prompt Questions in Detailed Reflection
MindfulDiary’s question-driven journaling process was also valued by participants as it assisted them with the process of daily reflections and documentation. Compared to their past experiences of journaling, where they had to reflect on their daily life by themselves, participants appreciated that MindfulDiary made the journaling process less daunting. P27 said, “Because I have to rely solely on my thoughts when I write alone, I sometimes get stuck. But when I was unsure about how to write, the AI helped me. I liked that part.”. The questions posed by MindfulDiary also guided participants in documenting their daily lives in a more detailed manner by asking their thoughts and feelings about a particular event. Such probing allowed for richer, more in-depth entries. P13 shared, “I used to write diaries on my own and just wrote whatever came to mind. MindfulDiary, however, helped me write in more detail because of the specific questions.”
6.3.4. Building Narratives: Structuring Daily Reflections with MindfulDiary
MindfulDiary’s capabilities, such as generating contextualized follow-up questions and summarizing the conversation, made the process easier for participants who struggled to organize daily thoughts and events underpsychotherapy (Donnelly and Murray, 1991). In their past experiences, our participants expressed difficulties in journaling because of disjointed thoughts, a lack of clarity in ideas, or inconsistencies in their stories. However, with the support of the LLM in the Mindful Diary, these challenges were addressed, motivating them to record their daily lives persistently. P13 remarked, ” I often had trouble putting sentences In the past, I would worry about writing the next part. But with this tool, I just tell the story of my day, and it seamlessly continues and wraps it up, presenting a well-structured diary entry. That’s its biggest advantage.”
6.4. MHPs’ Perspectives on MindfulDiary for Clinical Mental Health Settings
6.4.1. Expert Usage of MindfulDiary
In this section, we described how MHPs utilized the expert interface and what expectations and limitations they perceived. For this purpose, we interviewed five psychiatrists who were in charge of outpatient treatment at University Hospital.
All of the psychiatrists emphasized the critical value of an expert interface based on information recorded in the daily lives of patients. Specifically, E3 highlighted the program’s value in that MindfulDiary consistently aids in recording daily entries, allowing them to utilize more detailed patient data during outpatient visits. “Patients, with the support of AI, can logically continue their narratives, ensuring more dialogue than a typical (paper-based) diary. This definitely aids me in my consultations.” (E3). In this section, we present some examples of how MindfulDiary was integrated into clinical settings.
Integrating MindfulDiary into Clinical Workflow. First, we investigate how psychiatrists use MindfulDiary in their workflow. We have not provided guidance on how to use the clinician dashboard. However, we informed psychiatrists that the LLM-based analysis may not be accurate, and therefore, it is necessary to check the interaction log-data. The psychiatrists reviewed MindfulDiary data from their patients every morning during the deployment study, which they typically spent reviewing patient chart data they were going to see that day. Depending on the severity and the focal concerns of the patient, psychiatrists spent about 5-10 minutes per patient reviewing the MindfulDiary data. After observing trends primarily through PHQ-9 in the clinician dashboard first, psychiatrists read summaries about overall events and documented emotions. If there were spikes or drops in the PHQ-9 or events/emotions that needed further verification, they checked the actual interaction log.
Enhancing Understanding and Empathy toward Patients. Psychiatrists indicated that MindfulDiary helped them gain a deeper understanding and empathy about their patients. They perceived that MindfulDiary served as a questioner that could elicit more objective and genuine responses from patients. Psychiatrists appreciated that the LLM was able to pose questions that might be sensitive or burdensome for them to ask, such as patients’ negative perceptions of their parents. E4 said: “There are times when it’s challenging to counter a patient’s narrative or offer an opposing perspective. For example, if a patient speaks very negatively about their mother, and we ask, ‘Didn’t she treat you well when you were younger?’, the patient might react aggressively, thinking, ‘Why is the therapist taking my mother’s side?’ However, since the LLM is a machine, such concerns are minimized.”.
Insights from Everyday Perspectives Outside Clinical Visits. Psychiatrists valued that MindfulDiary provided them with an understanding of patients’ conditions that would be difficult to gain during outpatient visits. For instance, E1 appreciated that MindfulDiary provided them with insights into patients’ positive feelings and experiences, which is typically difficult to obtain during clinical consultations. “Usually, when patients come for a consultation, they talk about bad experiences. Few people come to psychiatry to say, ‘I’ve been doing well.’ Even if they have good things to say, they usually don’t bring them up. But I was happy to see that there were many positive statements in these records, like ’I did that and felt good.’ Especially in depression, the presence or absence of positive emotions is crucial. It’s a good sign if they show such positive responses.”. E2 envisioned its potential application to medication management, which is another critical aspect of psychiatric care. He thought these records could be used as a window into understanding how patients react to and perceive medications. For patients undergoing drug therapy, “If the primary treatment method is pills, but they don’t seem to have an effective response or there’s a decline in medication acceptance, I could potentially understand the reasons for it through this diary.” (E2).
Understanding Patient Progress Through Consistent Record-Keeping. Feedback from patients highlighted that interactions with MindfulDiary made it easier for patients to maintain a consistent record, as it mitigated the challenges associated with recording. Psychiatrists perceived that having consistent daily data offered them opportunities to observe trends in a patient’s condition. E2 said: “From our perspective as clinicians, even though we might only see a patient once a month, having access to a record of how they’ve been throughout the month would allow us to track their progress, which is highly beneficial.”. In particular, the ability to examine changes not only through quantitative tools like the PHQ-9 but also using a qualitative approach can offer a comprehensive understanding and shed light on the mechanisms influencing a patient’s mental health.
6.4.2. Perceived Opportunities and Limitations of MindfulDiary
While mental health MHPs generally appraised the utility of the MindfulDiary positively, they also perceived potential limitations, sharing considerations when integrating MindfulDiary into clinical settings.
Significance of Tone and Manner in Patient Data Analysis. Although patient data summarized and extracted in the expert interface effectively aided in understanding the patient, experts thought that the summarized texts would not convey the patient’s tone, pace, and other nuances, which are integral to the Mental Status Examination (MSE) that clinicians utilize. However, MHPs identified the opportunity to perform such analysis from the raw data that patients entered. As the MSE measures objective and quantitative aspects, incorporating such an analysis could make significant improvements in understanding the patient. E1 said, “In the same way as P14, understanding the tone of this patient may also be possible. That’s because we use something called psychiatric MSE, where we observe more than just the patient’s appearance, such as tone, pace, and more. Even a short analysis of one’s linguistic behavior would be great.”
Potential Misuses and Concerns around MindfulDiary. In our field study, one patient participant perceived the MindfulDiary as a channel to convey their intentions and situations to their psychiatrist. Specifically, the participant, P9, talked to their psychiatrist, “Have you seen what I wrote?”, which indicated that the patient was actively attempting to share their current state and situation through MindfulDiary. In spite of the fact that such usage did not seem problematic per se, one psychiatrist raised concerns about the possibility that patients with borderline personality disorders might misuse MindfulDiary as a weapon to manipulate others, such as their providers and parents. “In some cases, people self-harm out of genuine distress, but others do it to manipulate others, instilling guilt in them so they’ll do what they want. There are some patients who write about their distress with sincerity, while there are some who exaggerate their distress in order to get attention.” For patients exhibiting symptoms of schizophrenia or delusions, there was a concern that MindfulDiary’s feature of revisiting past entries could act as a feedback loop, developing and amplifying their delusions. E2 said, “This diary lets you revisit and organize your past actions. For schizophrenia patients with delusions or unique beliefs, referencing past writings might reinforce their pre-existing delusions. Reaffirming ’Yes, I’m right’ can be problematic. The LLM’s summaries could exacerbate these delusions if they emphasize distorted content.”
7. Discussion
In this study, we present MindfulDiary, an LLM-driven journal designed to document the daily experiences of psychiatric patients through naturalistic conversations. Here, we reflect on the opportunities presented by LLM-driven journaling for psychiatric patients and discuss considerations for integrating an LLM-driven patient system into the clinical setting.
7.1. Guiding patient journaling through conversations offering diverse perspectives
Our study highlighted the potential of MindfulDiary in clinical settings, mainly where adherence to interventions is important (Musiat et al., 2022). Core symptoms of depression, such as loss of energy, difficulty in carrying out mental processes, and feelings of apathy, often contribute to lower adherence to a professional’s advice or intervention (Kennedy, 2022). Clinicians who participated in our FGI also highlighted these challenges in motivating patients to utilize the diary writing program. Our findings demonstrated that MindfulDiary helped mitigate these challenges by transforming the conventional journaling process into engaging conversations. Using MindfulDiary, users were able to engage in conversations with the system by answering prompts and questions, which made them feel the journaling process was more accessible and intriguing. This active participation ensures that the users are not overwhelmed by the task and are guided in documenting their feelings and experiences more richly.
Depression often locks patients into negative and rigid thought patterns (Beck, 1979). Such patterns, resistance to change established thought paradigms, can severely limit a patient’s ability to perceive issues from multiple angles, leading to a harsh self-judgment (Marazziti et al., 2010). Our study highlighted that the varied perspectives offered by LLM-driven chatbots like MindfulDiary could help challenge such fixed viewpoints (Gero et al., 2022). By prompting users to revisit their initial evaluations or suggest alternative viewpoints, these chatbots could help break the cycle of cognitive rigidity. While our research underscores the promising role of LLM-driven chatbots in assisting psychiatric patients’ journaling process, it’s essential to note that these are preliminary findings. More work is needed to substantiate these findings in a clinical context.
7.2. Chatbot as a Mediator for Fostering Patient Understanding
Studies have suggested that sharing the data captured via chatbots with others, such as health professionals and family members, could further serve as an effective mediator that helps convey more truthful information (Lukoff et al., 2018; Lee et al., 2020a). For instance, patients consistently displayed deep self-disclosure through chatbots, whether or not they intended to share their inputs with health professionals (Lee et al., 2020a). Aligned with prior work on PGHD (Cohen et al., 2016; Nundy et al., 2014), MHPs in our study also perceived that MindfulDiary has shed light on patients’ daily events, emotions, and thoughts that might have been difficult to gain through regular clinical visits. This data offered MHPs valuable insights into the patient’s experiences and context. However, our study findings point to the need to carefully consider. In our study, we observed that the act of sharing journal content with MHPs through MindfulDiary led some patients to exaggerate their conditions or needs. This case underscores that when integrating a system like MindfulDiary into a clinical context, not only the design of chatbots that lead to patient disclosure behavior but also the complex patient-provider dynamics in clinical settings play a significant influence. The growing prevalence of chatbots in mental health domains emphasizes the need for a holistic approach to their design and implementation. We highlight that engineers and MHPs need to collaborate closely, ensuring that these tools are not only technically sound but also tailored to meet the intricate dynamics of clinical settings (Thieme et al., 2022).
7.3. Considerations for Integrating LLM-driven Patient Program
In this section, we discuss the consideration for integrating LLM-driven patient programs into clinical mental health settings, drawing insights from the design and evaluation of MindfulDiary.
Aligning domain experts’ expectations of LLMs. Developing and deploying MindfulDiary, we learned that aligning MHPs’ expectations with the capabilities and limitations of LLMs involves significant challenges. The capability of generative language models to improve mental health is difficult to measure in comparison with AI models in other medical domains, where objective metrics can determine performance. For instance, in medical imaging, AI can be evaluated based on its accuracy in identifying target diseases from MRI scans, using precise numerical percentages of correct identifications (Shen et al., 2017). On the other hand, in the realm of mental health chatbots, gauging success is more nuanced, as it involves subjective interpretations of emotional well-being and psychological improvement, which cannot be easily quantified or compared in the same straightforward manner. This challenge is amplified in mental health, where soft skills like rapport building and emotional observation are important (Gabbard and Crisp-Han, 2017). The use of LLMs in the mental health field is emerging, but little has been said about evaluating or defining the performance of models that are tailored to mental health. Our iterative evaluation process involving MHPs could inform researchers about how to develop and evaluate LLM-mediated mental health technology. When integrating into the clinical setting, this evaluation is also necessary for anticipating who the system would target and for what purpose it would be used. Hence, we advocate that engineers and researchers should carefully consider how to assist domain experts, who may lack AI expertise, in fully and accurately grasping the role and operation of LLM. It is also crucial for researchers and engineers to collaborate closely with these professionals to ensure the technology aligns with therapeutic needs and best practices (Thieme et al., 2022).
Tailored LLM Evaluation for Clinical Mental Health Domains. The domain of mental health, which our study addresses, is characterized by the vulnerability of its target user group. The content discussed within this domain is often emotionally charged and sensitive. Therefore, prioritizing user safety becomes even more essential in this domain than in others. Considering the sensitivity of the domain, during our evaluation process, MHPs thoroughly tested the LLM’s output by trying out conversations on various sensitive topics in both implicit and explicit ways, drawing upon their clinical experiences. The contents the MHPs input were much more diverse and wide-ranging than what engineers could generate during the development. Additionally, MHPs showed concern that the hallucinations of the language model could reinforce or expand the delusions of patients with delusional disorders. We highlight that developing evidence-based tests or benchmark sets to anticipate the behavior of the language models in collaboration with MHPs is critical when leveraging LLMs for clinical mental health settings.
Role of MHPs in Integration. Considering the limitations of current LLMs (Korngiebel and Mooney, 2021), it is critical to involve MHPs when deploying LLM-driven systems for patients in mental health contexts. While planning the field deployment study of MindfulDiary, we identified specific roles that MHPs could play. In the pre-use phase, MHPs should determine the suitability of users and facilitate the onboarding process with patients. During the mid-use phase, they should closely monitor interactions with the LLM and be prepared to intervene in cases of crises or unexpected use scenarios. Furthermore, they can offer or adjust treatments based on long-term data periodically. Additionally, they should regularly re-evaluate the continued use of the system. While some of these tasks should carefully be designed not to burden MHPs too much, it is important that LLMs do not make autonomous decisions about patients (e.g., diagnosis, prescription, or crisis management) but instead operate under professional oversight.
7.4. Limitations and Future Work
Our recruitment method could impact the generalizability of our findings, as we recruited the patient participants for our field study from a single university hospital. Although we aimed to recruit patients with diverse types and levels of symptoms, our participants are not representative samples of psychiatric patients. They were young (mostly adolescents) and consulted by a fixed number of psychiatrists. While this work is just a first step toward designing an LLM-driven journaling app for psychiatric patients, further investigation is necessary with subjects from various backgrounds.
To implement our pipeline, we used OpenAI’s GPT API, which provided the most capable LLM at the time of our study and was accessible via commercial API. As GPT models are continually updated, later models may not yield the same conversational behavior. To generalize the performance of our conversational pipeline design, future work is needed to compare multiple versions of MindfulDiary with different underlying LLMs.
8. Conclusion
In this paper, we designed MindfulDiary to assist psychiatric patients undergoing outpatient treatment with journaling in their daily lives. Keeping the clinical mental health setting in mind, our system was developed in collaboration with MHPs, from the initial concept building to the design of LLM’s conversation flow and evaluation. MindfulDiary leverages a stage-based LLM-driven chatbot, enabling patients to interact through prompt questions and answers, while complying with guidelines based on MHPs and literature. We conducted a field deployment study with 28 patients over 4 weeks. We found that the versatility, narrative-building capability, and diverse perspectives provided by MindfulDiary assisted patients in consistently enriching their daily records. The enriched records from MindfulDiary provided psychiatrists with deeper insights, enhancing their understanding and empathy toward their patients. We hope that this research provides a case study and insight into the development of an LLM-driven chatbot for mental health that is clinically relevant and reflects the needs and experiences of MHPs.
Acknowledgements.
We thank our study participants for their time and efforts. We also thank Eunkyung Jo and Yubin Choi for providing feedback on the early draft of this paper. This work was supported as a research internship at NAVER AI Lab.